# Unified Multi-task Learning for Voice-Based Detection of Diverse Clinical Conditions

### v1

@@ 4.2.4. Latent Space Visualization
To qualitatively assess the discriminative capacity of the representations learned by our model, we visualize the penultimate-layer embeddings (the 512-dimensional vector $\mathbf{z}$) using t-distributed Stochastic Neighbor Embedding (t-SNE) (Van der Maaten and Hinton, 2008). Figure 4 presents these visualizations for two representative tasks: Alzheimer’s Disease/MCI (left panel) and Parkinson’s Disease (right panel), comparing embeddings from positive (disease) cases against negative (control) cases.
@@ Appendix A SHAP-Based Feature Attribution for Acoustic Biomarkers
The objectives of this analysis were twofold: (1) to independently validate whether traditional, interpretable acoustic features are predictive of voice-related disorders, and (2) to identify which specific acoustic descriptors are most influential for each classification task using SHAP (SHapley Additive Explanations) (Lundberg and Lee, 2017).
@@ A.2. Identifying Salient Acoustic Features with SHAP
Clinical Interpretation of Feature Relevance. Notably, the acoustic features identified as salient by SHAP correspond to well-established physiological correlates. Perturbations in jitter and shimmer reflect vocal fold instability and irregular phonation, commonly associated with laryngeal or neurogenic dysphonia (Lopes et al., 2017). Another acoustic feature Spectral_gravity, which measures energy distribution across frequencies, is indicative of airflow limitation and has been linked to compromised respiratory mechanics (Taylor et al., 2020). Cepstral features such as cepstral_peak_prominence relate to harmonic structure and are sensitive to glottal closure behavior, which is relevant in vocal fold paralysis and tremor (Balasubramanium et al., 2011). Similarly, formant-based features like mean_F2_loc capture shifts in vocal tract resonance and articulatory posture, which are often impaired in neurodegenerative disorders (Mirarchi et al., 2017). Collectively, these associations reinforce the biological validity of acoustic biomarkers and highlight their value as non-invasive indicators of clinical conditions, thereby supporting the foundational premise of voice-based disease modeling adopted in our primary framework.

### v2

@@ 4.2.4. Latent Space Visualization
To qualitatively assess the discriminative capacity of the representations learned by our model, we visualize the penultimate-layer embeddings (the 512-dimensional vector $\mathbf{z}$) using t-distributed Stochastic Neighbor Embedding (t-SNE) (van2008visualizing). Figure 4 presents these visualizations for two representative tasks: Alzheimer’s Disease/MCI (left panel) and Parkinson’s Disease (right panel), comparing embeddings from positive (disease) cases against negative (control) cases.
@@ Appendix A SHAP-Based Feature Attribution for Acoustic Biomarkers
The objectives of this analysis were twofold: (1) to independently validate whether traditional, interpretable acoustic features are predictive of voice-related disorders, and (2) to identify which specific acoustic descriptors are most influential for each classification task using SHAP (SHapley Additive Explanations) (lundberg2017unified).
@@ A.2. Identifying Salient Acoustic Features with SHAP
Clinical Interpretation of Feature Relevance. Notably, the acoustic features identified as salient by SHAP correspond to well-established physiological correlates. Perturbations in jitter and shimmer reflect vocal fold instability and irregular phonation, commonly associated with laryngeal or neurogenic dysphonia (lopes2017accuracy). Another acoustic feature Spectral_gravity, which measures energy distribution across frequencies, is indicative of airflow limitation and has been linked to compromised respiratory mechanics (taylor2020age). Cepstral features such as cepstral_peak_prominence relate to harmonic structure and are sensitive to glottal closure behavior, which is relevant in vocal fold paralysis and tremor (balasubramanium2011cepstral). Similarly, formant-based features like mean_F2_loc capture shifts in vocal tract resonance and articulatory posture, which are often impaired in neurodegenerative disorders (mirarchi2017signal). Collectively, these associations reinforce the biological validity of acoustic biomarkers and highlight their value as non-invasive indicators of clinical conditions, thereby supporting the foundational premise of voice-based disease modeling adopted in our primary framework.